OpenStudy (anonymous):

Lisa has only nickels and dimes in her money box. She knows that she has less than $15 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. Which of the following statements best describes the steps to graph the solution to the inequality in x and y?

OpenStudy (anonymous):

Draw a dashed line to represent the graph of 5x + 10y = 1500, and shade the portion below the line for positive values of x and y. Draw a dashed line to represent the graph of 5x + 10y = 1500, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 1500, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x – 5y = 1500, and shade the portion below the line for positive values of x and y.

OpenStudy (anonymous):

I know the answer. It's eggplant. Gimme a medal.

OpenStudy (anonymous):

Dang you, Gakupo. @misssunshinexxoxo @alexistheking777

OpenStudy (misssunshinexxoxo):

Ignore her she will eventually get banned

OpenStudy (anonymous):

He's a guy. I know him and he's just trying to bug me.

OpenStudy (misssunshinexxoxo):

5x would accurately represent it, same with 10y for dimes. she has LESS than $15 in her box; shade downwards. If she wanted to have more than $15 or wants more than $15, it would be upward

OpenStudy (misssunshinexxoxo):

What do you feel the answer is by my explanation?

OpenStudy (misssunshinexxoxo):

Would like to make sure you understand

OpenStudy (anonymous):

Well, I don't really understand the "dashed line" thingy

OpenStudy (misssunshinexxoxo):

A would be the answer the dash line is for representation of graph

OpenStudy (anonymous):

Oh otay. Thank you again!

OpenStudy (anonymous):

Your welcome! :P I gave you a medal because I WUV U. I also sent you an eggplant. Check you doormat for a box of eggys. <3

OpenStudy (anonymous):

Seriously?

OpenStudy (misssunshinexxoxo):

Reporting kamui

OpenStudy (misssunshinexxoxo):

@Luigi0210

OpenStudy (misssunshinexxoxo):

@horsegirl27

OpenStudy (anonymous):

Nooooo I'm sorreh @misssunshinexxoxo